OpenStudy (otakulyfe):

Solve the system of equations. 2x + 3y = 6 x + y = 4

OpenStudy (otakulyfe):

@.Sam.

OpenStudy (.mehek.):

Would you like to solve using substitution or elimination?

OpenStudy (otakulyfe):

it doesn't say it just says solve

OpenStudy (benlindquist):

what do u want to use?

OpenStudy (.mehek.):

I know, so which method would like to use to solve?

OpenStudy (otakulyfe):

elimination

OpenStudy (.mehek.):

Multiply everything in the second equation by -2 to eliminate x -2(x + y = 4)

OpenStudy (otakulyfe):

so that would make it -2x+2y=-8

OpenStudy (.mehek.):

2x + 3y = 6 -2x - 2y = -8 You would add those equations together 2x + -2x = ? 3y + -2y = ? 6 + -8 = ?

OpenStudy (otakulyfe):

-2

OpenStudy (.mehek.):

for all of them?

OpenStudy (otakulyfe):

x 1y -2

OpenStudy (.mehek.):

2x + -2x = 0 so we eliminated x the others are correct so we have y = -2

OpenStudy (.mehek.):

So for the solution, we have (? , -2) Now we need to find x

OpenStudy (.mehek.):

To do that, you can chose any of the 2 equations and out -2 instead of y and solve for x

OpenStudy (otakulyfe):

2x+3(-2)=6

OpenStudy (.mehek.):

Yes good :) Now 3 * -2 = ?

OpenStudy (otakulyfe):

-6

OpenStudy (.mehek.):

Ok so we have 2x - 6 = 6 Can you solve for x?

OpenStudy (otakulyfe):

x=6 is that correct

OpenStudy (.mehek.):

Correct so for x we have 6 and y we have -2 The solution would be (6, -2)

OpenStudy (otakulyfe):

thank you so much for the help

OpenStudy (.mehek.):

You're welcome :)